OpenStudy (anonymous):

what is the standard form of the line y=-x+3/5? a) -5x-5y=3 b) 5x-5y=3 c) -5x+5y=3 d) 5x+5y=3

OpenStudy (angela210793):

D i guess

OpenStudy (radar):

The "general form" is ax + by +c = 0 For this case it would be: -x-y+3/5=0 or -5x-5y+3=0 or 5x+5y -3=0 or modified slightly 5x+5y=3

OpenStudy (anonymous):

D is the correct answer because. y=-x+3/5. Multiply everything by 5 to get rid of the fractions. 5y=-5x+3, then add 5x to each side. 5y+5x=3